A clown standing on a 23.5m high platform tosses ball vertically upward into the air at a speed of 6.5 m/s. The ball leaves his hand 3.4m above the platform. (A) how high from the floor will it go? (B) suppose the clown fqils to catch the ball, with what velocity will it hit the floor?


1
Expert's answer
2020-11-10T06:56:31-0500

a)


"H=H_0+h+\frac{v^2}{2g}=23.5+3.4+\frac{6.5^2}{2(9.8)}=29.0\ m"

b)


"V^2=v^2+2g(H_0+h)\\V^2=6.5^2+2(9.8)(23.5+3.4)\\V=23.9\frac{m}{s}"